# Two-color partitions with evens in one color

### v1

@@ Abstract.
integer partitions, two-color partitions, minimal excludant, $q$-series.
@@
First author partially supported by Simons Foundation Grant 633284
@@ 1. Introduction
$\sum_{n\geq 0}F(n)q^{n}=\frac{1}{(q;q^{2})_{\infty}^{2}(q^{2};q^{2})_{\infty}}=1+2q+4q^{2}+8q^{3}+14q^{4}+24q^{5}+\cdots.$ (1.1)
@@
$\sum_{n\geq 0}F_{0}(n)q^{n}=\frac{(q^{16},q^{6},q^{10};q^{16})_{\infty}}{(q;q^{2})_{\infty}(q^{2};q^{2})_{\infty}^{2}}.$
@@
For any nonnegative integer $n$, we have that $F_{0}(n)$ equals the number of partitions of $n$ in two colors (red and blue) wherein the odd parts and the parts $\equiv 6,10,16\ (\mathrm{mod}\ 16)$ may appear only in the blue color.
@@
$\sum_{n\geq 0}F_{1}(n)q^{n}=q\frac{(q^{16},q^{2},q^{14};q^{16})_{\infty}}{(q;q^{2})_{\infty}(q^{2};q^{2})_{\infty}^{2}}.$
@@
For any nonnegative integer $n$, we have that $F_{1}(n)$ equals the number of partitions of n1n-1 in two colors (red and blue) wherein the odd parts and the parts $\equiv 2,14,16\ (\mathrm{mod}\ 16)$ may appear only in the blue color.
@@
The confirms that $F_{2}(4)=p_{4,2}(4,\text{blue})$. Furthermore, we have $\overline{p}_{4,2}(4,\text{blue})=4$ counting
@@
$(-q;q)_{\infty}=\frac{1}{(q;q^{2})_{\infty}},$ (1.2)
@@
This rest of the paper is organized as follows. In Section 2, we establish more connections between our sequences and integer overpartitions. In Section 3 we provide the proof of Theorems 1-2. In Section 4 we give the proof of Theorems 3-4. Finally, Section 5 is devoted to some comments including new $q$-series identities and open problems which are related to our work in this paper.
@@ 2. Connections with overpartitions
$F_{0}(n)$ $=\frac{\overline{p}(n)+\overline{p}\big(\frac{n}{2}\big)}{2},$ (2.1)
$F_{1}(n)$ $=\frac{\overline{p}(n)-\overline{p}\big(\frac{n}{2}\big)}{2},.$ (2.2)
@@
It easily seen that
@@
$\sum_{n\geq 0}H(n)q^{n}=(-q;q^{2})_{\infty}^{2}(-q^{2};q^{2})_{\infty}=1+2q+2q^{2}+4q^{3}+6q^{4}+8q^{5}+\cdots.$ (2.5)
@@
$\sum_{n\geq 0}\big(H_{2}(n)-H_{3}(n)\big)q^{n}=(q^{2};q^{2})_{\infty}(q;q^{2})_{\infty}^{2}=1+2\sum_{n=1}^{\infty}(-1)^{n}q^{n^{2}}.$ (2.9)
@@ 3. Proof of Theorems 1-2
$=\frac{1}{(-q;q^{2})_{\infty}(q;q)_{\infty}},$ (3.1)
@@
$=\frac{1}{(q;q^{2})_{\infty}(q;q)_{\infty}}.$ (3.2)
@@
$=\frac{1}{2}\frac{1}{(q;q)_{\infty}(q^{2};q^{2})_{\infty}}\sum_{n=-\infty}^{\infty}q^{2n^{2}+n}\big(1+(-1)^{n}\big)$ (3.3)
@@
$=\frac{(q^{16},q^{6},q^{10};q^{16})_{\infty}}{(q;q)_{\infty}(q^{2};q^{2})_{\infty}}$ (3.4)
$=\frac{(q^{16},q^{6},q^{10};q^{16})_{\infty}}{(q;q^{2})_{\infty}(q^{2};q^{2})_{\infty}^{2}},$
@@
$=\frac{1}{2}\frac{1}{(q;q)_{\infty}(q^{2};q^{2})_{\infty}}\sum_{n=-\infty}^{\infty}q^{2n^{n}+n}\big(1-(-1)^{n}\big)$
@@
$=q\frac{(q^{16},q^{2},q^{14};q^{16})_{\infty}}{(q;q^{2})_{\infty}(q^{2};q^{2})_{\infty}^{2}}.$
This completes the proof.
@@ 4. Proof of Theorems 3-4
$(q,-xq,-1/x;q)_{\infty}=\sum_{n=-\infty}^{\infty}x^{n}q^{\frac{n(n+1)}{2}}.$ (4.1)
Proof of Theorem 3. It clear that
$\sum_{n\geq 0}\big(F_{2}(n)-F_{3}(n)\big)q^{n}$ $=\frac{1}{(-q^{2};q^{2})_{\infty}(q;q^{2})_{\infty}^{2}},$ (4.2)
$\sum_{n\geq 0}\big(F_{2}(n)+F_{3}(n)\big)q^{n}$ $=\frac{1}{(q^{2};q^{2})_{\infty}(q;q^{2})_{\infty}^{2}}.$ (4.3)
@@
$=\frac{1}{2(q^{2};q^{2})_{\infty}(q;q^{2})_{\infty}^{2}}\big(1+(q^{2};q^{2})_{\infty}(q^{2};q^{4})_{\infty}\big)$ (4.4)
@@
$=\frac{1}{2(q^{2};q^{2})_{\infty}(q;q^{2})_{\infty}^{2}}\Big(1+\sum_{n=-\infty}^{\infty}(-1)^{n}q^{2n^{2}}\Big)$ (4.5)
@@ 5. Concluding remarks
1. Our proofs for Corollaries 1-2 and for Theorems 3-4 are analytic relying of $q$-series manipulations.
@@
$+\sum_{n\geq 0}\frac{q^{2n+1}}{(q^{2};q^{2})_{n}(q;q^{2})_{n+1}(q;q^{2})_{n}}\quad(\text{greatest part odd-red}).$ (5.1)
@@
$+\sum_{n\geq 0}\frac{q^{2n+1}}{(q^{2n+2},q^{2n+1},q^{2n+1};q^{2})_{\infty}}\quad(\text{smallest part odd-red}).$ (5.2)
@@
$=\frac{1}{(q^{2};q^{2})_{\infty}(q;q^{2})_{\infty}^{2}}.$ (5.3)
@@
$+\sum_{n\geq 0}q^{2n+1}(-q^{2};q^{2})_{n}(-q;q^{2})_{n+1}(-q;q^{2})_{n}\quad(\text{greatest part odd-blue}).$ (5.4)
@@
$+\sum_{n\geq 0}q^{2n+1}(-q^{2n+2},-q^{2n+3},-q^{2n+3};q^{2})_{\infty}\quad(\text{smallest part odd-blue}).$ (5.5)

### v2

@@ Abstract.
integer partitions, two-color partitions, minimal excludant, $q$-series
@@
First author partially supported by Simons Foundation Grant 633284.
@@ 1. Introduction
(1.1) $\sum_{n\geq 0}F(n)q^{n}=\frac{1}{(q;q^{2})_{\infty}^{2}(q^{2};q^{2})_{\infty}}=1+2q+4q^{2}+8q^{3}+14q^{4}+24q^{5}+\cdots.$
@@
$\sum_{n\geq 0}F_{0}(n)q^{n}=\frac{(q^{16},-q^{6},-q^{10};q^{16})_{\infty}}{(q;q^{2})_{\infty}(q^{2};q^{2})_{\infty}^{2}}.$
@@
For any nonnegative integer $n$, we have that $F_{0}(n)$ equals the number of partitions of $n$ in two colors (red and blue), with optional overlines, wherein the odd parts and the parts $\equiv 0\ (\mathrm{mod}\ 16)$ may appear only in the blue color; the parts $\equiv 6,10\ (\mathrm{mod}\ 16)$ may appear in red and blue, with the first blue occurrence of each such part optionally overlined; and all other even parts may appear in either color.
@@
$\sum_{n\geq 0}F_{1}(n)q^{n}=q\frac{(q^{16},-q^{2},-q^{14};q^{16})_{\infty}}{(q;q^{2})_{\infty}(q^{2};q^{2})_{\infty}^{2}}.$
@@
For any nonnegative integer $n$, we have that $F_{1}(n)$ equals the number of partitions of n1n-1 in two colors (red and blue), with optional overlines, wherein the odd parts and the parts $\equiv 0\ (\mathrm{mod}\ 16)$ may appear only in the blue color; the parts $\equiv 2,14\ (\mathrm{mod}\ 16)$ may appear in red and blue, with the first blue occurrence of each such part optionally overlined; and all other even parts may appear in either color.
@@
This confirms that $F_{2}(4)=p_{4,2}(4,\text{blue})$. Furthermore, we have $\overline{p}_{4,2}(4,\text{blue})=4$ counting
@@
(1.2) $(-q;q)_{\infty}=\frac{1}{(q;q^{2})_{\infty}},$
@@
The rest of the paper is organized as follows. In Section 2, we establish more connections between our sequences and integer overpartitions. In Section 3 we provide the proof of Theorems 1-2. In Section 4 we give the proof of Theorems 3-4. Finally, Section 5 is devoted to some comments including new $q$-series identities and open problems which are related to our work in this paper.
@@ 2. Connections with overpartitions
(2.1) $F_{0}(n)$ $=\frac{\overline{p}(n)+\overline{p}\big(\frac{n}{2}\big)}{2},$
(2.2) $F_{1}(n)$ $=\frac{\overline{p}(n)-\overline{p}\big(\frac{n}{2}\big)}{2},.$
@@
It is easily seen that
@@
(2.5) $\sum_{n\geq 0}H(n)q^{n}=(-q;q^{2})_{\infty}^{2}(-q^{2};q^{2})_{\infty}=1+2q+2q^{2}+4q^{3}+6q^{4}+8q^{5}+\cdots.$
@@
(2.9) $\sum_{n\geq 0}\big(H_{2}(n)-H_{3}(n)\big)q^{n}=(q^{2};q^{2})_{\infty}(q;q^{2})_{\infty}^{2}=1+2\sum_{n=1}^{\infty}(-1)^{n}q^{n^{2}}.$
@@ 3. Proof of Theorems 1-2
(3.1) $=\frac{1}{(-q;q^{2})_{\infty}(q;q)_{\infty}},$
@@
(3.2) $=\frac{1}{(q;q^{2})_{\infty}(q;q)_{\infty}}.$
@@
(3.3) $=\frac{1}{2}\frac{1}{(q;q)_{\infty}(q^{2};q^{2})_{\infty}}\sum_{n=-\infty}^{\infty}q^{2n^{2}+n}\big(1+(-1)^{n}\big)$
@@
(3.4) $=\frac{(q^{16},-q^{6},-q^{10};q^{16})_{\infty}}{(q;q)_{\infty}(q^{2};q^{2})_{\infty}}$
$=\frac{(q^{16},-q^{6},-q^{10};q^{16})_{\infty}}{(q;q^{2})_{\infty}(q^{2};q^{2})_{\infty}^{2}},$
@@
$=\frac{1}{2}\frac{1}{(q;q)_{\infty}(q^{2};q^{2})_{\infty}}\sum_{n=-\infty}^{\infty}q^{2n^{2}+n}\big(1-(-1)^{n}\big)$
@@
$=q\frac{(q^{16},-q^{2},-q^{14};q^{16})_{\infty}}{(q;q^{2})_{\infty}(q^{2};q^{2})_{\infty}^{2}}.$
In the last step we applied (4.1) with $q\to q^{16}$ and $x=q^{-2}$. This completes the proof.
@@ 4. Proof of Theorems 3-4
(4.1) $(q,-xq,-1/x;q)_{\infty}=\sum_{n=-\infty}^{\infty}x^{n}q^{\frac{n(n+1)}{2}}.$
Proof of Theorem 3. It is clear that
(4.2) $\sum_{n\geq 0}\big(F_{2}(n)-F_{3}(n)\big)q^{n}$ $=\frac{1}{(-q^{2};q^{2})_{\infty}(q;q^{2})_{\infty}^{2}},$
(4.3) $\sum_{n\geq 0}\big(F_{2}(n)+F_{3}(n)\big)q^{n}$ $=\frac{1}{(q^{2};q^{2})_{\infty}(q;q^{2})_{\infty}^{2}}.$
@@
(4.4) $=\frac{1}{2(q^{2};q^{2})_{\infty}(q;q^{2})_{\infty}^{2}}\big(1+(q^{2};q^{2})_{\infty}(q^{2};q^{4})_{\infty}\big)$
@@
(4.5) $=\frac{1}{2(q^{2};q^{2})_{\infty}(q;q^{2})_{\infty}^{2}}\Big(1+\sum_{n=-\infty}^{\infty}(-1)^{n}q^{2n^{2}}\Big)$
@@ 5. Concluding remarks
1. Our proofs for Corollaries 1-2 and for Theorems 3-4 are analytic, relying on $q$-series manipulations.
@@
(5.1) $+\sum_{n\geq 0}\frac{q^{2n+1}}{(q^{2};q^{2})_{n}(q;q^{2})_{n+1}(q;q^{2})_{n}}\quad(\text{greatest part odd-red}).$
@@
(5.2) $+\sum_{n\geq 0}\frac{q^{2n+1}}{(q^{2n+2},q^{2n+1},q^{2n+1};q^{2})_{\infty}}\quad(\text{smallest part odd-red}).$
@@
(5.3) $=\frac{1}{(q^{2};q^{2})_{\infty}(q;q^{2})_{\infty}^{2}}.$
@@
(5.4) $+\sum_{n\geq 0}q^{2n+1}(-q^{2};q^{2})_{n}(-q;q^{2})_{n+1}(-q;q^{2})_{n}\quad(\text{greatest part odd-blue}).$
@@
(5.5) $+\sum_{n\geq 0}q^{2n+1}(-q^{2n+2},-q^{2n+3},-q^{2n+3};q^{2})_{\infty}\quad(\text{smallest part odd-blue}).$